OpenStudy (anonymous):

Solve the following system of equations by substitution. x + y = 5 y = 3 A. (-2, 3) B. (2, -3) C. (0, 0) D. (2, 3)

OpenStudy (anonymous):

These are solved by having an equation that describes the value of one variable and substitute that in the other equation. Since you already have the value of y, just substitute that in the first equation, what do you get? x+3=2 so what's x?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Crap! I apologize. I didn't read it correctly! The equation is x+y=5 So x+3=5. My bad!

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

Right! and since you know y=3, your solutions are (2,3)

OpenStudy (anonymous):

The answer is D. Substitute y=3 in the first equation and you would get x=2. Therefore, the solution is (2,3)

OpenStudy (anonymous):

thank you